Find the discriminant for each of the following quadratic equations and use it to determine the number of roots or real solutions it has.
5. x2 - 2x + 8 = 0
6. -x2+5x+6= 0
7. x2 - 14x + 49 = 0

Answer by venugopalramana(3286)

Find the discriminant for each of the following quadratic equations and use it to determine the number of roots or real solutions it has.
5. x2 - 2x + 8 = 0
FOR THE EQN.
AX^2+BX+C=0
D= B^2-4AC IS THE DISCRIMINANT
D=(-2)^2-4(1)(8)=4-32=-28...-VE
HENCE 2 DISTINCT IMAGINARY ROOTS
6. -x2+5x+6= 0
D=25+24=49..+VE & PERFECT SQUARE
HENCE 2 DISTINCT REAL RATIONAL ROOTS
7. x2 - 14x + 49 = 0
D=196-4*49-0
HENCE REAL EQUAL ROOTS.ONE ROO